スペクトル・散乱理論とその周辺

研究集会

Spectral and Scattering Theory and Related Fields

京都大学数理解析研究所の共同研究事業の一つとして、下記のように研究集会を催しますので、ご案内申し上げます。

研究代表者山田修宣（立命館大理工）

記

日時： ２０００年１２月４日（月）１３：３０から１２月６日（水）１５：４０まで場所： 京都大学数理解析研究所４階４２０号室京都市左京区北白川追分町市バス 農学部前 または 北白川 下車

プログラム

12 月 4 日（月）

13:30–14;10 中村 周 Shu Nakamura （東大数理）

L^{p}-Theory of spectral shift function and the Wigner estimate 14:20–15:00 吉冨 和志 Kazushi Yoshitomi （九大数理）

Eigenvalue problems on domains with cracks 15:20–16:00 勝田 篤 Atsushi Katsuda （岡山大理）

BC-method and stability of Gel’fand inverse spectral problem 16:10–16:50 望月 清 Kiyoshi Mochizuki （都立大理）

Jäger–Rejto approach on growth estimates

12 月 5 日（火）

9:30–10:10 門脇 光輝 Mitsuteru Kadowaki （都立航空高専）

On existence of scattering solutions for dissipative systems 10:20–11:00 広島 文生 Fumio Hiroshima （摂南大）

Gibbs measures and a boson localization of a scalar quantum field model

11:10–11:50 Ingrid Beltita (Inst. Math., Romanian Academy)

On an abstract radiation condition 13:30–14:10 Arne Jensen (Dept. Math. Sci., Aalborg University)

A unified approach to resolvent expansions at thresholds 14:20–15:00 足立 匡義 Tadayoshi Adachi （神戸大理）

Scattering theory for N-body quantum systems in a time-periodic electric field

15:20–16:00 磯崎 洋 Hiroshi Isozaki （阪大理） Asymptotic properties of solutions to 3–particle Schrödinger equations

16:10–16:50 一瀬 孝 Takashi Ichinose （金沢大理）

田村 英男 Hideo Tamura （岡山大理） On the norm convergence of the Trotter–Kato product formula with error bound

12 月 6 日（水）

9:30–10:10 中西 賢次 Kenji Nakanishi （神戸大理） Nonrelativistic limit of scattering theory for nonlinear Klein-Gordon equations

10:20–11:00 谷島 賢二 Kenji Yajima （東大数理）

張 果平 Gouping Zhang （東大数理） Smoothing property of Schrödinger equations with potentials superquadratic at infinity

11:10–11:50 伊藤 宏 Hiroshi Ito (愛媛大工）

田村 英男 Hideo Tamura （岡山大理） Scattering by magnetic fields at large separation in two dimensions

13:20–14:00 一ノ瀬 弥 Wataru Ichinose （信州大理） The Feymann path integral representation of Green functions of position and momentum operators

14:10–14:50 峯 拓矢 Takuya Mine （京大理） On negative eigenvalues of Schrödinger operators in magnetic fields with Robin boundary conditions

15:00–15:40 大鍛冶 隆司 Takashi Ōkaji （京大理） Absence of eigenvalues of Maxwell operators